OpenStudy (anonymous):

Estimate the value of 17^(1/4) using differentials??!!???

OpenStudy (anonymous):

Please show the steps

OpenStudy (anonymous):

Use google.

OpenStudy (anonymous):

I google it for you, hold on.

OpenStudy (anonymous):

actually I can't sorry, I have no idea what differentials are.

ganeshie8 (ganeshie8):

say \(\large y = x^{1/4}\) and suppose your starting value is x = 16

ganeshie8 (ganeshie8):

then dx = 1, yes ?

ganeshie8 (ganeshie8):

\(\large y = x^{1/4} \implies dy = \frac{1}{4}x^{-3/4} dx\)

ganeshie8 (ganeshie8):

plugin x = 16 and dx = 1 above ^^

OpenStudy (anonymous):

ganeshie8. why does dx=1? I don't see how you got that value

ganeshie8 (ganeshie8):

we have assumed our starting value is x=16 because 16^(1/4) = 2, which is easy to compute dx = 17-16 = 1

ganeshie8 (ganeshie8):

dx is the difference between final value of x and the starting value

OpenStudy (anonymous):

THANKS!